OpenStudy (anonymous):

Find points of inflection and discuss the concavity of the graph of the function. F(x)=2x^3-3x^2-12x+5

OpenStudy (anonymous):

I have so far F ' (x) = \[F'(x) = 6x^2-6x-12\] which gives you the critical #s of 2,-1 then \[F''(x) = 12x-6\] so possible inflection point is \[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

you've got the inflexcion point. now, test for concavity

OpenStudy (anonymous):

check the sign of the first derivative for any value of "x"

OpenStudy (anonymous):

ok so I do not need to know the x intercept or the y intercept.

OpenStudy (anonymous):

no wait.. there is no point of inflexion!!!!

OpenStudy (anonymous):

\[ F''(2)=12(2)-6>0\\ F''(-1)=12(-1)-6<0 \]

OpenStudy (anonymous):